OpenStudy (vera_ewing):

How can (1/3)x – 2 = y and (1/4)x + 11 = y be set up as a system of equations? A. 3y – x = –6 4y – x = 44 B. 3y + x = –6 4y + x = 44 C. 3y – 3x = –6 4y – 4x = 44 D. 3y + 3x = –6 4y + 4x = 44

OpenStudy (vera_ewing):

I think it's C

OpenStudy (unklerhaukus):

hmm, are they the only options?

OpenStudy (vera_ewing):

yes those are the only options. i tried to get help yesterday, but nincompoop said all of the choices were wrong so I'm confused :/

OpenStudy (unklerhaukus):

\(\tfrac13x-2=y\] multiplying the first equation by three \[x-6=3y\] rearranging \[x-3y=6\] multiply by negative one, to get a match with one of the equations in the options

OpenStudy (vera_ewing):

So it must be A?

OpenStudy (unklerhaukus):

similarly, for the second equation \[\tfrac14 x + 11 = y \] multiply by the denominator (four in this case) and rearrange to get the constants to the right-hand-side

OpenStudy (vera_ewing):

So the answer is A?

OpenStudy (unklerhaukus):

yep

OpenStudy (vera_ewing):

thanks! can you help me with a couple more?